# Local-Cloud Inference Offloading for LLMs in Multi-Modal, Multi-Task, Multi-Dialogue Settings

### v1

@@
Abstract
Compared to traditional machine learning models, recent large language models (LLMs) can exhibit multi-task-solving capabilities through multiple dialogues and multi-modal data sources. These unique characteristics of LLMs, beyond their large size, make their deployment more challenging during the inference stage. Specifically, (i) deploying LLMs on local devices faces computational, memory, and energy resource issues, while (ii) deploying them in the cloud cannot guarantee real-time service and incurs communication/usage costs. In this paper, we design a local-cloud LLM inference offloading (LCIO) system, featuring (i) a lightweight local LLM that can process simple tasks at high speed and (ii) a large-scale cloud LLM that can handle multi-modal data sources. LCIO employs resource-constrained reinforcement learning (RCRL) to determine where to make the inference (i.e., local vs. cloud) and which multi-modal data sources to use for each dialogue/task, aiming to maximize the long-term reward (which incorporates response quality, latency, and usage cost) while adhering to resource constraints. We also propose M4A1, a new dataset that accounts for multi-modal, multi-task, multi-dialogue, and multi-LLM characteristics, to investigate the capabilities of LLMs in various practical scenarios. We demonstrate the effectiveness of LCIO compared to baselines, showing significant savings in latency and cost while achieving satisfactory response quality.
Index Terms:
@@ Abstract
I Introduction
Large language models (LLMs) have demonstrated remarkable general-purpose task-solving capabilities across various fields, garnering significant global interest, and gradually integrating into daily life. Currently, renowned LLMs from industry giants include OpenAI’s GPT-4 [1], Google’s Gemma [2, 3] and Gemini [4], and Meta’s Llama [5, 6], among others. These industry giants have trained different LLM architectures on various datasets, exhibiting distinct performance levels and generalization capabilities.
Figure 1: Application Scenario for the LCIO System: Kitchen Activity Assistance with LLM. Upon receiving the user’s text prompt, RL determines the most suitable LLM (i.e., local LLM vs. cloud LLM) and the most relevant data modalities based on the task and previous dialogues. For the specific query, “Where did I put my dishes?” RL identifies that this question is most closely related to the first-person view and overhead view modalities. Consequently, the corresponding prompt and the selected modalities are uploaded to the cloud LLM, which processes the multi-modal inputs to respond to the query. The local LLM, with its limited capability (e.g., processing only text data), may not be able to solve this task.
@@ I Introduction
Motivation. The development of LLMs is thriving, yet deploying LLMs on local devices presents practical challenges, including computational, memory, and energy resources [7, 8, 9, 10, 11]. These hinder the installation of large-scale LLMs locally or real-time, seamless communication with cloud LLMs. These issues hinder the deployment of large-scale LLMs on users’ local devices [12, 13]. Industry giants have developed strategies that either (i) involve lightweight versions of LLMs, such as Google’s Gemma and Meta’s Llama, which offer solutions with varying numbers of parameters (typically 3B, 8B, 13B, etc.) for fast and cost-efficient response, or (ii) place LLMs in the cloud, where users interact with them via the internet, as seen with the ChatGPT application on mobile phones. The former often suffers from inferior inference performance, while the latter requires substantial network bandwidth and incurs high costs from service providers. Therefore, our motivation is to build a system that combines the advantages of both to enhance response quality while ensuring lower latency and usage costs.
Challenges. However, different from traditional machine learning models, achieving this objective with LLMs poses new challenges. LLMs operate in complex environments characterized by (i) multi-modal data sources (i.e., the LLM may need to process different types of data such as text or image to make reliable decisions), (ii) multi-task (i.e., users may want to solve different types of tasks such as message editing or recommendation), and (iii) multi-dialogue (i.e., users engage in multiple interactions with the LLM). Although these elements enrich the capabilities of LLMs, they introduce significant challenges in the inference stage. First, it is challenging to decide the optimal location for inference, i.e., whether to process by the local LLM or to offload tasks to the cloud LLM. This decision critically impacts not only the response quality but also the latency and usage costs. For example, local LLM inference may limit response quality due to the lightweight model deployed on local devices, whereas cloud LLM inference, although potentially more powerful, can introduce significant communication/computation delay [14]. Second, selecting which multi-modal data sources to use for each task/dialogue is non-trivial. Although various types of input data can improve the response quality of LLMs, not all tasks require comprehensive multi-modal data. For example, a message editing task may not require images as inputs, whereas some human activity tasks require various types of image modalities [15, 16, 17]. Deciding which modalities are necessary and how to efficiently process them without incurring excessive delay or cost is a critical aspect of deploying LLMs. Third, due to the expensive nature of LLM inference, online training strategies (i.e., real-time interaction with LLM inference) are often impractical. Consequently, it becomes necessary to generate a dataset for offline training. However, the inherent uncertainty of LLM inference presents additional challenges, as even identical inputs can produce different outputs. This uncertainty significantly exacerbates the difficulty of evaluating LLM outputs, particularly in multi-task or generative tasks where no standard answers or evaluation metrics exist.
Motivated by these, we aim to answer the following question:
In the local-cloud LLM system, (i) how to select the appropriate LLM for inference (i.e., local vs. cloud), (ii) how to choose the most informative and associated multi-modal data sources for each task and dialogue, while taking into account response quality, latency, and usage cost, and (iii) how to address the inherent uncertainty in LLM inference during offline training?
Contribution. We propose a local-cloud inference offloading (LCIO) strategy within a local-cloud LLMs system aimed at enhancing response quality, and reducing latency and usage costs. Fig. 1 illustrates the workflow of LCIO. We address the above research question by developing reinforcement learning (RL) that learns optimal actions in the current state through offline learning of the environment in which users interact with LLMs. The state is determined by the historical information of LLMs and modalities, and the current task (which reflects the task’s difficulty and relevance to the environment). The action involves selecting the LLM and the modalities to answer the aforementioned key research question. Our main contributions can be summarized as follows  [^1] The dataset and code is available at https://github.com/liangqiyuan/LCIO.:
  • We design a local-cloud LLM inference offloading system, termed LCIO, to optimize LLMs’ response quality, latency, and usage costs. We implement the LCIO system within a multi-modal, multi-task, multi-dialogue scenario. Specifically, we explore the capabilities of LLMs across various tasks, including assistant, query, recommendation, and message editing, and demonstrate its effectiveness in dynamically adapting to diverse conversational demands (Sec. III-A and Sec. III-B).
  • We propose resource-constrained RL (RCRL) to maximize long-term reward when utilizing the LCIO system. By selecting the best LLM and modalities for inference, this approach allows for effective trade-offs between response quality, latency, and usage costs. In addition, the reward takes into account the association between user’s text prompts and other modalities to quantify the connections between tasks and multi-modal data sources, incorporating them into the decision-making process (Sec. III-C).
@@
  • We introduce a nearest neighbor strategy for estimating the response quality of uncertain state-action pairs. This approach addresses two types of uncertainty: the non-deterministic evaluation (NDE), where identical state-action pairs can yield varying response quality evaluations, and the out-of-distribution (OOD) scenario, where the system must estimate the response quality for previously unseen state-action pairs. For the first scenario, our strategy provides an estimation of the response quality to reduce the uncertainty introduced by the LLM-as-judge. For the second scenario, it quantifies the potential responses that LLMs may generate when exploring any action in RCRL (Sec. III-D).
  • We propose and generate a new dataset termed M4A1, which considers multi-modal, multi-task, multi-dialogue, and multi-LLM characteristics, encapsulating these four “multi” elements all in one dataset. It includes (i) three different view images, (ii) four distinct tasks, (iii) two to five sequential dialogues, and (iv) four LLMs for different purposes. Based on this dataset, we explore the impacts of various RL algorithms, trade-off strategies, resource constraints, local devices, and response quality differences between LLMs to demonstrate the efficacy and versatility of the proposed LCIO system (Sec. IV).
Organization. The remainder of this paper is organized as follows. Sec. II reviews the related work. Sec. III presents our proposed LCIO system along with the design of RCRL. In Sec. IV, we introduce our M4A1 dataset and detail the experimental results. Finally, Sec. V concludes the paper and discusses future research directions.
II Related Work
II-A Efficient LLM on Local Devices
The rapid development of LLMs has profoundly influenced how people interact with AI, yet it also poses multiple challenges related to operational costs, energy consumption, and environmental impact. To mitigate inference overhead, researchers have proposed various solutions. For example, OpenAI has optimized every layer of the stack and introduced an end-to-end multi-modal LLM, GPT-4o, which achieves twice the inference speed and halves the inference cost [18]. Chen et al.[19] have explored strategies such as shrinking prompt sizes, linking multiple queries to share prompts, caching and reusing previous responses, fine-tuning cheaper LLMs with outputs from more expensive ones, and selectively mixing various LLM APIs. Beyond these architectural optimizations, various LLM architectures suitable for local devices [12, 20] and quantization techniques [21] have been proposed to enable deployment in resource-constrained devices.
Other edge-based LLM technologies, such as Apple Intelligence [22], have demonstrated remarkable capabilities on mobile devices, including the integration of ChatGPT into Siri. However, this solution still relies solely on user control to choose which LLM to use, rather than autonomously determining the most appropriate LLM or modality. This limitation is particularly evident when dealing with multiple multi-modal data sources. For example, consider an iPhone equipped with a front camera, a rear camera, and the ability to capture screenshots. The user must manually decide which data source aligns best with the current task (e.g., switching cameras) before uploading the selected content to ChatGPT, thereby introducing inconvenience and operational overhead. Moreover, existing frameworks (i.e., LLM selection strategies) fail to account for the intricate relationships between various modalities and tasks, nor do they consider the context or dialogue history across multiple turns.
II-B LLM Deployment over Network
LLMs over network, such as the ChatGPT application on mobile phones using cellular networks or WiFi, bring the convenience of queries anytime and anywhere to people’s lives, even facilitating multi-modal queries by accessing the phone’s camera [23, 24]. However, this also imposes requirements on the phone’s network connectivity and involves costs for LLM services. Such LLMs over the network present practical challenges for applications that require low latency [25]. For example, [26] discuss using GPT-4 for voice-assisted autonomous driving. The LLM response time is around 1.5 seconds, and there is also the risk of losing connection. Therefore, using LLM over the network can only serve as an assistive tool and is far from sufficient for autonomous driving. Some researchers have proposed solutions to reduce the latency of LLMs over networks. [27] proposed using LLM to generate latency-sensitive code paths during an initial offline phase prior to task execution, rather than during the system’s online operation, to reduce communication latency with cloud LLMs. [28] introduced TREACLE, an RL algorithm that selects different LLMs and prompts based on the inference costs and latency associated with each LLM. However, most existing frameworks either focus solely on one type of resource or only on the accuracy of LLM responding to simple inquiries, without considering the trade-offs and optimization of multiple key resource indicators in multi-modal, multi-task, multi-dialogue scenarios.
II-C Model and Modality Selection
Model and modality selection has been optimized on the basis of a range of criteria such as performance, cost, and latency. This is achieved through decision-making strategies that incorporate objective optimization, simulated annealing, genetic algorithms, among others [29]. Entropy has been used to establish a threshold to decide whether to offload inference tasks to a server [30]. Furthermore, a multi-objective optimization approach has been proposed, which accounts for the Shapley value of different modalities, communication overhead, and the recency of modality selection [31].
Although the aforementioned methods focus primarily on meeting specific performance standards or finding global optimal solutions, they are not designed to consider maximization of cumulative rewards throughout the decision-making process, which is crucial in time-series tasks. [32] developed RL for modality selection in a series of medical diagnostics. This method determines whether to collect additional modality data, which modality to collect, and when to conclude the examination and make a diagnosis, thereby maximizing the trade-off between accuracy and cost. However, this approach does not account for multi-disease diagnostic scenarios, nor does it consider the association between diagnostic content (i.e., disease) and modalities. [33] employed RL to realize resource allocation in LLM inference offloading, yet the approach does not consider multi-modal data or the time-series aspects of multi-dialogues. This omission fails to align with the practical scenarios of using LLMs to solve complex tasks, which typically involve interactions over extended periods and require the processing of diverse types of data. Differently from prior works, we implement these concepts in a multi-dialogue LLM scenario, not only maximizing cumulative rewards, but also considering the relevance between tasks and modalities in a multi-task setting to select the optimal models and modalities.
III Proposed LCIO System
III-A Multi-Task/Dialogue Scenario with Multi-Modal Data
@@
In our problem setting, a user aims to solve multiple types of task through a series of dialogues with the LLM. Each dialogue $t$ involves (i) a text prompt $P_{t}$, such as “Please recommend two dishes that I can cook based on what I have,” and (ii) a set of other modalities $\mathcal{M}$, such as different views of images, 3D points, or other formats of data describing the current situation. During the conversation with the LLM, a user may ask additional questions using prompts potentially related to the previous dialogues: “Where did I put my dishes?” or “Please turn on all the lights in the kitchen” (see Fig. 2). Here, each text prompt belongs to one of the $N$ different task categories, such as assistant, recommendation, query, and message editing. Moreover, distinct tasks have varying levels of difficulty and require different types of multi-modal information to solve. For example, the prompt “Where did I put my dishes?” requires additional modalities (e.g., images) for the LLM to solve the task, while the prompt “Please turn on all the lights in the kitchen” may not require other sources of data, indicating that the task is simpler. These characteristics within a multi-task, multi-dialogue scenario with multi-modal data sources distinguish our setup from many existing works that consider simpler settings (e.g., a single dialogue with text only), motivating the need for a new solution.
Figure 2: Challenges Emerging in the Local-Cloud LLM Inference. When the user poses questions of varying complexity and scenario information needs, it becomes challenging to determine which LLM and modality(s) should be used for inference. First-person view captures the specific action of the user and the objects but does not reveal the exact location of the cabinet. Overhead view provides information about the entire scenario but lacks details about what the person is specifically doing. Side view, similar to the overhead view, offers scenario information that overlaps in terms of the specific user action.
III-B Local-Cloud LLM Inference
Due to the inherent limitations of local devices in terms of computational resources, memory, and energy consumption, deploying large-scale LLM directly on these devices is often impractical. On the other hand, relying exclusively on cloud-based LLMs can introduce latency and potential usage costs, especially when handling multi-modal data. We propose the local-cloud LLM inference offloading (LCIO) system, as shown in Fig. 1, which strategically distributes tasks between local and cloud platforms to exploit their respective strengths. It consists of two different LLMs: the local LLM and the cloud LLM. The local LLM is a lightweight LLM (e.g., Phi-3-mini [34]) deployed on local devices to handle simple tasks efficiently. This model is optimized for high-speed text prompt processing, suitable for tasks such as message editing, setting timers, or adjusting air conditioning settings, where a sophisticated understanding of multi-modal data is not critical. On the other hand, the cloud LLM is a large-scale multi-modal LLM (e.g., GPT-4o [18]) situated in the cloud capable of managing complex task and multi-modal data that exceed the processing capabilities of local devices. This LLM can integrate multi-modal data input, thus providing richer, more accurate responses where required. The advantage of the LCIO approach is its flexibility and efficiency. In the LCIO system, the user can flexibly choose to either offload inference to the cloud or process it locally, depending on the task’s difficulty and current resource requirements. By offloading computationally intensive tasks to the cloud, we preserve local resources while still benefiting from the cloud’s powerful computational abilities. This configuration allows each part of the system to operate within its capacity, optimizing overall performance and user experience (e.g., latency and costs).
Formally, we describe the interaction with either the local or cloud LLM at each dialogue $t$ as follows:
$R_{t}=\begin{cases}\text{LLM}_{\text{Local}}\left(P_{t}\right),\ \ \ \text{forlocal},\\ \text{LLM}_{\text{Cloud}}\left(P_{t}\right),\ \ \ \text{for cloud text-only},\\ \text{LLM}_{\text{Cloud}}\Big{(}P_{t},\bigcup\limits_{m\in\mathcal{M}_{t}}\{D^{(m)}\}\Big{)},\\ \ \ \ \ \ \ \ \ \ \text{for cloud with modalities in $\mathcal{M}_{t}$},\end{cases}$ (1)

### v3

@@
Abstract.
Compared to traditional machine learning models, recent large language models (LLMs) can exhibit multi-task-solving capabilities through multiple dialogues and multi-modal data sources. These unique characteristics of LLMs, together with their large model size, make their deployment more challenging. Specifically, (i) deploying LLMs on local devices faces computational, memory, and energy resource issues, while (ii) deploying them in the cloud cannot guarantee real-time service and incurs communication/usage costs. In this paper, we design TMO, a local-cloud LLM inference system with Three-M Offloading: Multi-modal, Multi-task, and Multi-dialogue. TMO incorporates (i) a lightweight local LLM that can process simple tasks at high speed and (ii) a large-scale cloud LLM that can handle multi-modal data sources. We develop a resource-constrained reinforcement learning (RCRL) strategy for TMO that optimizes the inference location (i.e., local vs. cloud) and multi-modal data sources to use for each task/dialogue, aiming to maximize the long-term reward (response quality, latency, and usage cost) while adhering to resource constraints. We also contribute M4A1, a new dataset we curated that contains reward and cost metrics across multiple modality, task, dialogue, and LLM configurations, enabling evaluation of offloading decisions. We demonstrate the effectiveness of TMO compared to several exploration-decision and LLM-as-Agent baselines, showing significant improvements in latency, cost, and response quality.
@@ Abstract.
journalyear: 2025copyright: ccconference: International Symposium on Theory, Algorithmic Foundations, and Protocol Design for Mobile Networks and Mobile Computing; October 27–30, 2025; Houston, TX, USAbooktitle: International Symposium on Theory, Algorithmic Foundations, and Protocol Design for Mobile Networks and Mobile Computing (MobiHoc ’25), October 27–30, 2025, Houston, TX, USAdoi: 10.1145/3704413.3764429isbn: 979-8-4007-1353-8/25/10ccs: Computing methodologies Distributed artificial intelligence
1. Introduction
@@
Large language models (LLMs) have demonstrated remarkable general-purpose task-solving capabilities across various fields and have been gradually incorporated into daily life. The development of LLMs is thriving, yet deploying LLMs on local devices presents practical challenges, including computational, memory, and energy resource limitations (Su et al., 2024; Wang et al., 2024). These issues hinder the deployment of large-scale LLMs on users’ local devices. Industry giants have developed strategies that either (i) involve lightweight versions of LLMs, such as Google’s Gemma and Meta’s LLaMA, which offer solutions with varying numbers of parameters (typically 3B, 8B, 13B, etc.) for fast and cost-efficient response, or (ii) place LLMs in the cloud, where users interact with them via the Internet, as seen with the ChatGPT application on mobile phones. The former often suffers from inferior inference performance, while the latter requires substantial network bandwidth and incurs high costs from service providers. This creates a pressing need to build systems that combine the advantages of both strategies, i.e., enhancing response quality while ensuring lower latency and usage costs.
Figure 1. Application scenario of the TMO system as an LLM Assistant. In this example, cloud LLM is selected along with first-person and overhead views, as the query requires visual information.
1.1. Research Challenges
Optimizing LLMs over networks introduces unique challenges beyond traditional machine learning models, as LLMs operate with (i) multi-modal data sources (e.g., text, images), (ii) multi-task environments (e.g., editing, recommendations), and (iii) multi-dialogue contexts (i.e., ongoing conversations with users). Although these elements enrich the capabilities of LLMs, they introduce significant challenges, particularly in the inference stage:
1. Determining the optimal inference location: Determining whether to process tasks using a local LLM or to offload them to a cloud LLM is non-trivial. This decision critically impacts not only the response quality but also the latency and costs. For example, local LLM inference may limit response quality due to the lightweight model deployed on local devices, whereas cloud LLM inference, although more powerful and capable of multi-modal processing, can introduce significant communication/computation delay.
2. Understanding relationships between modalities, tasks, and dialogues: There are intricate associations between multi-modal data sources, inference tasks, and their temporal variation across multiple dialogues. Although various types of input data can improve the response quality of LLMs, not all tasks require comprehensive multi-modal data. For example, a message editing task may not require images as inputs, whereas some human activity tasks require various types of image modalities (Lee et al., 2024). Furthermore, when considering multi-dialogue scenarios, where each dialogue represents a task in a sequential order, earlier uploaded data modalities may provide sustained informational advantages, yet they might necessitate re-uploading if the information becomes outdated.
3. Uncertainty issues in offline training with LLM inference data: It is often desirable to train auxiliary models using LLM inference results, e.g., to learn modality-task associations. The computationally expensive nature of LLMs renders it impractical to conduct online training of such models. Consequently, generating datasets for offline training becomes necessary. However, due to inherent uncertainty in LLM inference, which arises from the probabilistic nature of language model outputs (Hu et al., 2023), such datasets may contain multiple identical samples with varying response scores. This uncertainty substantially complicates the evaluation of LLM outputs, particularly in multi-task or generative scenarios where standard answers or evaluation metrics do not exist.
Motivated by these challenges, we aim to answer the following research questions. In the context of local-cloud LLM systems:
  • (RQ1)
    How can we optimally select the appropriate LLM for inference (i.e., local LLM vs. multi-modal cloud LLM)?
  • (RQ2)
    How can we identify and utilize the most informative and relevant multi-modal data when necessary for tasks in dialogues, while balancing response quality, latency, and usage cost?
  • (RQ3)
    How can we effectively address the inherent uncertainty in LLM inference data during offline training?
1.2. Summary of Contributions
In this paper, we develop TMO, a local-cloud LLM inference offloading optimization system shown in Fig. 1. TMO orchestrates “Three-M” Offloading – multi-modal, multi-task, and multi-dialogue – aimed at enhancing response quality and reducing latency and usage costs. We address the above research questions through a resource-constrained reinforcement learning (RCRL) methodology that learns optimal actions (i.e., RQ1 and RQ2) to respond to mutli-task and multi-dialogue scenarios. We develop an estimator of LLM inference through a nearest neighbor strategy for effective offline learning (i.e., RQ3). Our main contributions are as follows:
  • We construct TMO to account for adaptations between device-side LLM and cloud LLM inference within multi-modal, multi-task, multi-dialogue settings. This includes formalizing the interaction with both types of LLMs to dynamically adapt with diverse conversational demands (Sec. 3.1 and Sec. 3.2).
  • We formulate TMO’s decision engine in terms of resource-constrained reinforcement learning (RCRL) to maximize cumulative reward across multi-dialogue interactions by selecting the optimal LLMs and modalities for inference. This approach enables a concrete optimization of trade-offs between response quality, latency, and usage costs while incorporating task-modality associations into the decision-making process (Sec. 3.3).
@@ 1.2. Summary of Contributions
  • We devise a nearest neighbor strategy to overcome state-action pair uncertainty in estimating response quality during offline RL training. Our approach addresses two types of uncertainty that manifest in LLM evaluation contexts: non-deterministic evaluation (NDE), where identical state-action pairs can yield varying response quality evaluations, and out-of-distribution (OOD), where the system estimates the response quality for unseen state-action pairs (Sec. 3.4).
  • We curate a new dataset, termed M4A1, which contains samples from three multi-modal data sources, four inference tasks, 2-5 dialogues, and four LLMs, with real-world measurements of latency and usage costs. Through extensive evaluation with this dataset, we demonstrate TMO’s superior performance in response quality, latency, and usage cost compared to two SOTA exploration-decision and 15 LLM-as-Agent baselines (Sec. 4).
2. Related Work
Efficient On-Device and Over-Network LLM. The rapid development of LLMs has profoundly influenced human-AI interaction while introducing significant challenges related to operational costs, energy consumption, and environmental impact. Researchers propose various solutions to enable deployment in resource-constrained devices (Chen et al., 2023), such as architectural optimization (Chu et al., 2023) and quantization (Lin et al., 2024). Furthermore, network-based approaches have emerged, such as (Shen et al., 2024) proposed the client-edge co-inference method, which partitions LLM into lightweight layers running on devices and parameter-heavy layers operating at the edge to enhance inference efficiency.
However, existing frameworks primarily focus on either single-modal data sources or single-query accuracy (Zhang et al., 2024), neglecting the complex trade-off optimization of key resource indicators (e.g., response quality, latency, costs) (Dong et al., 2024) and failing to address the intricate relationships between multi-modal, multi-task, and multi-dialogue context (i.e., RQ1 and RQ2). Furthermore, (Zhang et al., 2024; He et al., 2024) employ datasets with fixed evaluation metrics, which may not fully represent the complexity in many real-world LLM applications. Moreover, online training methods incur prohibitively expensive costs and lead to irreproducible resource consumption, particularly when considering human subjects or alternative approaches for evaluating LLM inference (i.e., RQ3). In contrast to previous work, we propose an offline training approach to preserve LLM inference results. While this inevitably introduces uncertainties in the LLM inference data, we address this challenge by further incorporating an uncertainty estimator for response scores.
Inference Selection and Offloading Optimization. More generally, there is a rich set of literature on strategies for optimizing local-edge-cloud inference location selection/offloading, based on metrics such as performance, cost, and latency (Singhal et al., 2024). These works explore diverse methodological approaches that can be categorized into network optimization strategies, such as confidence-based device-server selection (Al-Atat et al., 2024) and energy-delay considerations for LLM training (Liu and Zhao, 2024), as well as RL frameworks, such as accuracy-delay-fairness device-server selection (Beytur et al., 2024) and multi-device offloading with quality-latency-energy considerations (Chen et al., 2024).
However, existing hierarchical inference approaches (Al-Atat et al., 2024; Beytur et al., 2024), which orchestrate offloading between local devices and cloud servers, only consider model size differences. They neglect fundamental distinctions in their input modalities, for example, one scenario is device-side LLMs being limited to text-only inputs, while server-side models can process multimodal inputs (i.e., RQ1). Furthermore, recall that the unique challenge in our settings is multi-dialogue interactions, yet these methods primarily focus on meeting specific performance standards or finding immediate optimal solutions without considering cumulative reward maximization throughout the decision-making process. (He et al., 2024) employed active inference for LLM inference offloading but disregarded multi-modal data and mutli-task in multi-dialogues, failing to align with practical LLM scenarios (i.e., RQ2). Differently from prior works, we implement these concepts in a multi-dialogue LLM scenario, not only maximizing cumulative rewards, but also considering the modality-task associations to select the optimal LLM and multi-modal data sources.
3. Proposed TMO System
3.1. Multi-Modal, Multi-Task, Multi-Dialogue Interaction
We consider a setup in which a user is aiming to solve multiple types of inference tasks through a series of dialogues with LLMs. Each dialogue $t$ involves:
  • a text prompt $P_{t}$ which describes the user’s intent, e.g., “Recommend dishes that I can cook with my ingredients,” and
  • a set of other modalities $\mathcal{M}_{t}\subseteq\mathcal{M}$, where $\mathcal{M}$ represents all modalities available to the user (e.g., different image views, 3D points, or other data formats describing the current situation).
@@ 3.1. Multi-Modal, Multi-Task, Multi-Dialogue Interaction
During the conversation with LLMs, a user may ask additional questions using prompts potentially related to the previous dialogues (e.g., “Where did I put my dishes?” or “Please turn on all the lights in the kitchen”). Here, we assume each text prompt belongs to one of $N$ different task categories (e.g., assistant, recommendation, query, and message editing). Moreover, distinct tasks have varying levels of difficulty and require different types of multi-modal information to solve. For example, the prompt “Where did I put my dishes?” requires additional modalities (e.g., images) for the LLM to solve the task, while the prompt “Please turn on all the lights in the kitchen” may not require other sources of data, indicating that the task is simpler. These characteristics within a multi-task, multi-dialogue scenario with multi-modal data sources distinguish our setup from existing works (Zhang et al., 2024; He et al., 2024; Yang et al., 2024) that consider simpler settings (e.g., a single dialogue with text only), motivating the need for a new solution.
3.2. Local-Cloud LLM Inference
As shown in Fig. 1, a key objective of TMO is to strategically distribute tasks between local and cloud platforms, to exploit their respective strengths. TMO consisiders two different LLMs:
  • The local LLM is a lightweight LLM (e.g., Phi-3-mini) deployed on local devices, intended to handle simple tasks efficiently and process text prompt understanding with high speed.
  • The cloud LLM is a large-scale multi-modal LLM (e.g., GPT-4o) situated in the cloud capable of managing complex tasks and processing information that exceeds local device capabilities, integrating multi-modal data input to provide richer and more accurate responses where required.
TMO determine whether to process LLM inference locally or offload to the cloud, depending on the task difficulty, available modalities, and resource requirements. Offloading computationally intensive tasks to the cloud enables preservation of local resources while still benefiting from the cloud powerful computational abilities.
Formally, we represent the user-LLM interaction at each dialogue $t$ in terms of three possibilities:
[…]